OpenStudy (anonymous):

what is the integral of ((x-3)/5) dx step by step

OpenStudy (freckles):

you can seperate the fraction (x-3)/5 is equal to 1/5*x-3/5

OpenStudy (freckles):

then separate the integral \[\int\limits \frac{1}{5} x dx - \int\limits \frac{3}{5} dx\] do you know how to continue ?

OpenStudy (anonymous):

Thanks @freckles